# Cavity Control of Topological Qubits: Fusion Rule, Anyon Braiding and Majorana-Schrödinger Cat States

### v1

@@ Abstract
Fusion rule.— MZMs are Ising anyons obeying the fusion rule $\sigma\times\sigma=1+\Psi$, where two Ising anyons ($\sigma$) fuse into either a vacuum ($1$) or a standard Dirac fermion ($\Psi$[21, 31]. Fusion rule experiments are in general simpler than braiding experiments, thus they are first-ever candidates for testing the non-Abelian features of MZMs.
@@
In the present cavity setup, for testing the non-trivial fusion rule only a pair of MZMs is required. We consider the fusion rule in the $1CK$ case by adiabatically increasing the KC-cavity coupling strength. Thus, we start with a single uncoupled KC in its ground state featuring a pair of MZMs located at the chain edges. During the whole process, the total fermionic parity of the KC states is fixed to be even ($+1$) as we start with the ground state of a non-interacting KC-cavity, i.e. $|\chi_{1}\rangle=|-\rangle_{z}\otimes|0_{\rm ph}\rangle$ where $|0_{\rm ph}\rangle$ denotes the photon vacuum state. After the cavity is strongly coupled to the KC two new MZMs emerge at the neighbor sites of the cavity as previously demonstrated in Ref. [25].
@@
One immediate consequence of the $H_{1}$ structure in Eq. (1) is the expression for its USC limit eigenstates: if $\lambda\gtrsim\omega\gg\Delta$, the matter-light eigenstates become uncoupled, given by the separable states $|\Psi\rangle\approx|-\rangle_{x}\otimes|\alpha\rangle$ where $\hat{\sigma}_{x}|-\rangle_{x}=-|-\rangle_{x}$ and the photon state is nothing but the coherent state $|\alpha\rangle$ with $\alpha=-\lambda/\omega$. Thus, the final matter state, after adiabatically turning on the KC-cavity coupling until the USC regime is reached, turns out to be $|-\rangle_{x}=\left(\lvert\circ\circ\rangle-|\bullet\bullet\rangle\right)/\sqrt{2}$ indicating once more that the cavity coupling doesn’t alter the total fermionic parity. Therefore, the final matter state is an equal probability superposition of states with no fermions on each side of the cavity and states with a single fermion on each side. The fusion outcome becomes probabilistic, not deterministic, as clearly indicated by the resulting KC-cavity state. This result corresponds to a positive test of the fusion protocol of two Ising anyons. In the case of quantum dot-based Kitaev systems the readout of the fusion outcome can be detected by standard capacitance measurements or tunneling signatures [32].
@@
Figure 3: Fusion rule protocol in a $1CK$ system. In panel (a), we depict the ground state fermion correlation function $P(s-1,s+1)$, with the dashed blue line representing the resonance condition. In panel (b), for the resonant case, we consider a time-dependent linear ramping of the radiation-matter interaction coupling strength, $\lambda=vt=t/t_{a}$ with $v=t_{a}^{-1}$ the ramping rate, where the evolution time $t\in\left[0,4t_{a}\right]$.
We numerically test the validity of the fusion protocol in $1CK$ system by diabatically turning on the matter-light coupling strength. For that purpose we evaluate the fermion correlation function of the two nearest neighbor Majorana fermions to the cavity, $P(s-1,s+1)=i\langle\gamma_{s-1,2},\gamma_{s+1,1}\rangle=-\langle\hat{\sigma}_{x}\rangle$. Figure 3(a) shows a 3D plot of that correlation in the ground state as a function of photon energy and coupling strength ($\phi=0$). To test for diabatic errors, we show the same correlation function in Fig. 3(b) for time-dependent cases with different velocities of turning on of the matter-photon coupling strength (different ramping rates). Notably, we obtain an excellent agreement between analytical and numerical results, even well beyond the realm of applicability of the USC regime, for sufficiently slow ramping processes.
@@
Braiding protocol.— As a further step, we propose a cavity-chain setup to perform an MZM braiding/exchange process. Most recipes for simulations of braiding protocols in 1D topological systems rely on either Floquet dynamics (Majorana time crystals [33]) or quantum dot mediated exchange [21, 34]. The former case demands extreme timing control while the latter scheme requires a vanishing dot energy level splitting which hampers an adiabatic evolution harming the goal of a geometric phase $\pi/4$, as required for a perfect exchange between two MZMs. Our goal is to reach the same $\pi/4$-phase by performing an adiabatic cyclic process on the cavity itself, i.e. on the coherent state $|\alpha\rangle$, by leveraging the separable property of the ground state (GS) in the USC regime, i.e. $\lvert{\rm GS}\rangle\simeq\lvert\chi\rangle\otimes|\alpha\rangle$ with $\lvert\chi\rangle$ a matter KC state. In our setup, the polariton KC-cavity state can be coupled to the Majorana state, and so is not involved in braiding, only initialization (and possibly readout).
Consider a $2CK$ acting on chain sites $j=s$ (coupling strength $\lambda_{1}e^{i\phi_{1}}$) and $j=s+1$ (coupling strength $\lambda_{2}e^{i\phi_{2}}$), governed by $H_{2}$ in Eq. (2). Let us start by fixing $\lambda_{1}>0$ in the USC regime and $\phi_{1}=0$, without any loss of generality, while $\lambda_{2}$ and $\phi_{2}$ can be tunable. The MF just to the left of the cavity, i.e. $\hat{\gamma}_{L}=\hat{\gamma}_{s-1,2}$, always behaves as a free MZM [25] while $\hat{\gamma}_{R}=\hat{\gamma}_{s+1,1}$ acquires the MZM condition only in the $\lambda_{2}\rightarrow 0$ regime. We will be focusing on the braiding process between these two MFs, $\hat{\gamma}_{L}$ and $\hat{\gamma}_{R}$.
For $\phi_{1}=0$, the composite ground state in the USC regime becomes
$\lvert{\rm GS}_{\pm}\rangle\simeq\begin{cases}-i\lvert\Psi_{+}\rangle\otimes|\alpha_{+}\rangle,&\text{for }\phi_{2}=0,\\ +i\lvert\Psi_{-}\rangle\otimes|\alpha_{-}\rangle,&\text{for }\phi_{2}=\pi.\end{cases}$ (3)
@@
Where $\lvert\Psi_{\pm}\rangle=(\lvert\bullet_{L,R},\bullet_{c_{1}},\circ_{s+1}\rangle\pm\lvert\circ_{L,R},\bullet_{c_{1}},\bullet_{s+1}\rangle)/\sqrt{2}$, $\alpha_{+}=-(\lambda_{1}+\lambda_{2})/(\sqrt{2}\omega)$, and $\alpha_{-}=-\lambda_{1}/(\sqrt{2}\omega)$ regardless the value of $\lambda_{2}\leq\lambda_{1}$. In the 1CK limit, i.e. $\lambda_{2}\rightarrow 0$, the system’s ground state turns out to be
$\begin{split}\lvert{\rm GS}\rangle&\simeq\frac{1}{\sqrt{2}}\left(\lvert{\rm GS}_{+}\rangle-\lvert{\rm GS}_{-}\rangle\right),\\ &=\lvert\bullet_{L,R},\bullet_{c},\circ_{s+1}\rangle\otimes|\alpha\rangle.\end{split}$ (4)
@@
With $\alpha=-\lambda_{1}/(\sqrt{2}\omega)$. The braiding or exchange process between $\hat{\gamma}_{L}$ and $\hat{\gamma}_{R}$ MFs, as described by the operator $\hat{U}_{B}^{(L,R)}=e^{\frac{\pi}{4}\hat{\gamma}_{L}\hat{\gamma}_{R}}=\frac{1}{\sqrt{2}}\left(1+\hat{\gamma}_{L}\hat{\gamma}_{R}\right)$ (performing the exchange transformations $\hat{\gamma}_{L}\rightarrow\hat{\gamma}_{R}$ and $\hat{\gamma}_{R}\rightarrow-\hat{\gamma}_{L}$), when acting on the ground state produces the result $\hat{U}_{B}^{(L,R)}\lvert{\rm GS}\rangle=e^{-i\frac{\pi}{4}}\lvert{\rm GS}\rangle$. By using a local phase shifter acting just on the $j=s+1$ chain site [35, 36, 37, 30], a controlled complex coupling strength $\lambda_{2}e^{i\phi_{2}}$ should be possible enabling a cyclic evolution of the cavity coherent state $|\alpha\rangle$ and consequently to access to a specific Berry phase of the whole matter-photon state (dynamic phases which depend on the system’s specific time evolution may be experimentally canceled out [38, 39]).
@@
Now, we focus on recovering the same $e^{-i\frac{\pi}{4}}$ global phase factor on the GS by performing an adiabatic cyclic evolution of the coherent cavity field (Berry phase). The GS adiabatic evolution is fully contained in the two-dimensional sub-space spanned by the states $\{\lvert{\rm GS}_{+}\rangle,\lvert{\rm GS}_{-}\rangle\}$, as previously assured (for details see the SM). We perform a cyclic evolution of the GS by fixing a $\lambda_{2}>0$ value while varying the phase $\phi_{2}=\phi$ of the coupling strength, $0\leq\phi\leq 2\pi$. Thus, for the GS we have $\lvert{\rm GS}(\phi)\rangle=C_{+}(\phi)\lvert{\rm GS}_{+}(\phi)\rangle+C_{-}(\phi)\lvert{\rm GS}_{-}\rangle$. Notice that the state $\lvert{\rm GS}_{-}\rangle$ does not depend on the phase $\phi$. We proceed with an analytical calculation of the Berry phase using the sub-space-confined evolution of the GS. By direct calculation, we obtain:
$\varphi_{B}=\int_{0}^{2\pi}\frac{d\phi}{2\omega}\delta E_{\phi}\left[1-\frac{\delta E_{\phi}}{\sqrt{4\Delta^{2}\lvert\langle\alpha_{-}\lvert\alpha_{+}(\phi)\rangle\lvert^{2}-\delta E^{2}_{\phi}}}\right].$ (5)
@@
The parity $P_{L,R}=i\langle\hat{\gamma}_{L}\hat{\gamma}_{R}\rangle$, between $\hat{\gamma}_{L}$ and $\hat{\gamma}_{R}$ MFs, is plotted in Fig. 4(b) as a function of $\lambda_{2}e^{i\phi_{2}}$, for $0\leq\lambda_{2}\leq\lambda_{1}$ and $0\leq\phi_{2}\leq 2\pi$. $P_{L,R}\rightarrow-1$ for $\lambda_{2}\approx 0$ in agreement with the GS given by Eq. (4). The solid red line shows the corresponding results for varying real $\lambda_{2}$ values from the positive to the negative region. At the same time, the closed curve depicts the $P_{L,R}$ evolution just at the $\lambda_{2}$ value producing the correct Berry phase of $\pi/4$, simulating the perfect braiding/exchange between $\hat{\gamma}_{L}$ and $\hat{\gamma}_{R}$ MFs.

### v2

@@ 3 Fusion rule
MZMs are Ising anyons obeying the fusion rule $\sigma\times\sigma=1+\Psi$, where two Ising anyons ($\sigma$) fuse into either a vacuum ($1$) or a standard Dirac fermion ($\Psi$[44, 54]. Fusion rule experiments are in general simpler than braiding experiments, thus they are first-ever candidates for testing the non-Abelian features of MZMs.
@@
To test the non-trivial fusion rule, we focus on the single-site cavity configuration (see Sec. 2.1), where only a pair of MZMs is required. Specifically, we implement the fusion protocol in the $1CK$ setup by adiabatically increasing the coupling strength between the Kitaev chain and the cavity. Thus, we start with a single uncoupled KC in its ground state featuring a pair of MZMs located at the chain edges. During the whole process, the total fermionic parity of the KC states is fixed to be even (+1+1) as we start with the ground state of a non-interacting KC-cavity, i.e. $|\chi_{1}\rangle=|-\rangle_{z}\otimes|0_{\rm ph}\rangle$ where $|0_{\rm ph}\rangle$ denotes the photon vacuum state. For a KC-cavity strongly coupled system two new MZMs emerge at the neighbor sites of the cavity.
@@
One immediate consequence of the $H_{1}$ structure in Eq. (4) is the expression for its USC limit eigenstates: if $\lambda\gtrsim\omega\gg\Delta$, the matter-light eigenstates become uncoupled, given by the separable states $|\Psi\rangle\approx|-\rangle_{x}\otimes|\alpha\rangle$ where $\hat{\sigma}_{x}|-\rangle_{x}=-|-\rangle_{x}$ and the photon state is nothing but the coherent state $|\alpha\rangle$ with $\alpha=-\lambda/\omega$. Thus, the final matter state, after adiabatically turning on the KC-cavity coupling until the USC regime is reached, turns out to be $|-\rangle_{x}=\left(|\circ\circ\rangle-|\bullet\bullet\rangle\right)/\sqrt{2}$ indicating once more that the cavity coupling doesn’t alter the total fermionic parity. Therefore, the final matter state is an equal probability superposition of states with no fermions on each side of the cavity and states with a single fermion on each side. The fusion outcome becomes probabilistic, not deterministic, as clearly indicated by the resulting KC-cavity state. This result corresponds to a positive test of the fusion protocol of two Ising anyons. In the case of quantum dot-based Kitaev systems the readout of the fusion outcome can be detected by standard capacitance measurements or tunneling signatures [55].
@@
Figure 3: Fusion rule protocol in a $1CK$ system. Panel (a) shows the ground-state fermionic correlation function P(s1,s+1)P(s-1,s+1), where the dashed blue line indicates the resonance condition. In panel (b), we consider the resonant case with a time-dependent linear ramp of the light-matter coupling strength, given by $\lambda(t)=\lambda_{\max}(t/t_{a})$, where $t_{a}$ is the annealing time and $\lambda_{\max}=4$ is the maximum coupling.
In typical experimental platforms, measuring the fusion rule involves evaluating the two-point correlation function or, equivalently, the probabilistic fermion parity [20, 44]. When a single nanowire initially in the topological phase with even fermion parity is split into two halves, the fermion parity readout between the resulting sectors is expected to yield values of $\pm 1$ with a zero average. This outcome reflects the parity combinations allowed by the fusion rules—either even-even or odd-odd—each occurring with an equal probability of $1/2$. To test the validity of the fusion protocol in the $1CK$ system, we numerically simulate the process by diabatically activating the matter-light coupling. Specifically, we compute the fermion correlation function between the two Majorana fermions nearest to the cavity, $P(s-1,s+1)=i\langle\gamma_{s-1,2},\gamma_{s+1,1}\rangle=-\langle\hat{\sigma}_{x}\rangle$. This analysis provides a quantitative means to verify the expected behavior of the fusion protocol under the given conditions. Figure 3(a) shows a 3D plot of that correlation in the ground state as a function of photon energy and coupling strength ($\phi=0$).
3.1 Time-dependent fusion rule control
In this subsection, we describe the numerical methods used to obtain the time-dependent results. The light-matter coupling is modeled by a linear ramp $\lambda(t)=\lambda_{\max}\,t/t_{a}$, where $t_{a}$ is the annealing time, controlling the ramping speed, and $\lambda_{\max}$ is the maximum coupling strength. At $t=0$, the cavity is decoupled from the Kitaev chain, and the system is initialized in the ground state of the Hamiltonian given in Eq. (4) with $\lambda=0$. As the coupling is gradually turned on, the two subsystems begin to interact, governed by the controlled annealing protocol.
[…]
The correlation function P(s1,s+1)P(s{-}1,s{+}1) serves as a key signature of the fermionic fusion rule. In Fig. 3(b), we present its behavior as a function of the annealing time, which corresponds to the duration over which the matter-photon coupling is gradually turned on. To assess the robustness of the protocol against diabatic errors, we compare this correlation for different ramping rates. Remarkably, we find excellent agreement between analytical predictions and numerical simulations, even for ramping protocols that extend beyond the strict boundaries of the ultrastrong coupling regime—provided the coupling is increased sufficiently slowly.
In summary, what a fusion rule experiment measures is, essentially, a two-point correlation or, equivalently, a probabilistic fermion parity [20, 44]. When a single nanowire, initially in the topological phase with even fermion parity, is split into two, the fermion parity readout between the two new sectors is expected to yield values of $\pm 1$ with zero average. This outcome reflects the allowed parity combinations, either even-even or odd-odd, with identical probabilities, $1/2$, and this is precisely what is here found for the $1CK$ case in the USC regime.
4 Braiding protocol
@@
As a further step, we propose a cavity-chain setup to perform an MZM braiding/exchange process. Most recipes for simulations of braiding protocols in 1D topological systems rely on either Floquet dynamics (Majorana time crystals [56]) or quantum dot mediated exchange [44, 57]. The former case demands extreme timing control while the latter scheme requires a vanishing dot energy level splitting which hampers an adiabatic evolution harming the goal of a geometric phase $\pi/4$, as required for a perfect exchange between two MZMs. Our goal is to reach the same $\pi/4$-phase by performing an adiabatic cyclic process on the cavity itself, i.e. on the coherent state $|\alpha\rangle$, by leveraging the separable property of the ground state (GS) in the USC regime, i.e. $|{\rm GS}\rangle\simeq|\chi\rangle\otimes|\alpha\rangle$ with $|\chi\rangle$ a matter KC state. In our setup, the polariton KC-cavity state can be coupled to the Majorana state, and so is not involved in braiding, only initialization (and possibly readout).
To simulate the braiding of two Majorana fermions, $\hat{\gamma}_{L}$ and $\hat{\gamma}_{R}$, we consider the $2CK$ setup in which the cavity couples to two adjacent sites in the bulk of the chain: $j=s$ and $j=s+1$. The corresponding light-matter interaction is governed by $H_{2}$ in Eq. (9), with site-dependent couplings $\lambda_{1}e^{i\phi_{1}}$ and $\lambda_{2}e^{i\phi_{2}}$, respectively. Without loss of generality, we fix $\lambda_{1}>0$ in the ultrastrong coupling regime and set $\phi_{1}=0$, while treating $\lambda_{2}$ and $\phi_{2}$ as tunable parameters.
 In this configuration, the Majorana fermion just to the left of the cavity, $\hat{\gamma}_{L}=\hat{\gamma}_{s-1,2}$, consistently behaves as a free MZM. In contrast, the Majorana to the right of the cavity, $\hat{\gamma}_{R}=\hat{\gamma}_{s+1,1}$, acquires the MZM character only in the limit $\lambda_{2}\rightarrow 0$. The braiding protocol we analyze involves these two MFs, $\hat{\gamma}_{L}$ and $\hat{\gamma}_{R}$, and is designed to explore how their exchange is mediated through cavity coupling dynamics.
For clarity, we schematically indicate which Majorana mode is being manipulated or exchanged at each stage of the protocol using dashed ovals in the inset of Fig. 4. This visual aid helps track the braiding steps and reinforces the interpretation of the cavity as a local control element enabling effective non-Abelian operations within a one-dimensional geometry.
As in the single-site cavity case, the weak-coupling regime of the two-site setup is best described using the matter states defined in Eq. (8), which offer a suitable basis for analyzing the dynamics governed by $\hat{H}_{2}$. However, in the USC regime, a more appropriate representation is given by the occupation number basis $\{|n_{L,R},n_{c1},n_{c2}\rangle\}$, where $f_{L,R}=\frac{1}{2}\left(\hat{\gamma}_{L}+i\hat{\gamma}_{R}\right)$ defines a non-local fermion composed of Majorana modes outside the cavity, and $c_{1}=\frac{1}{2}\left(\hat{\gamma}_{s,1}+i\hat{\gamma}_{s,2}\right)$ and $c_{2}=\frac{1}{2}\left(\hat{\gamma}_{s+1,1}+i\hat{\gamma}_{s+1,2}\right)$ are two local fermions fully contained within the cavity region.
Further details on this basis transformation and its physical implications are provided in A. For $\phi_{1}=0$, the composite ground state in the USC regime becomes
$|{\rm GS}_{\pm}\rangle\simeq\Bigg{\{}\begin{array}[]{cc}-i|\Psi_{+}\rangle\otimes|\alpha_{+}\rangle,&{\rm for}\;\phi_{2}=0,\\ +i|\Psi_{-}\rangle\otimes|\alpha_{-}\rangle,&{\rm for}\;\phi_{2}=\pi.\end{array}$ (11)
@@
Where $|\Psi_{\pm}\rangle=(|\bullet_{L,R},\bullet_{c_{1}},\circ_{s+1}\rangle\pm|\circ_{L,R},\bullet_{c_{1}},\bullet_{s+1}\rangle)/\sqrt{2}$, $\alpha_{+}=-(\lambda_{1}+\lambda_{2})/(\sqrt{2}\omega)$, and $\alpha_{-}=-\lambda_{1}/(\sqrt{2}\omega)$ regardless the value of $\lambda_{2}\leq\lambda_{1}$. In the 1CK limit, i.e. $\lambda_{2}\rightarrow 0$, the system’s ground state turns out to be
$|{\rm GS}\rangle$ $\simeq\frac{1}{\sqrt{2}}\left(|{\rm GS}_{+}\rangle-|{\rm GS}_{-}\rangle\right),$ (12)
$=|\bullet_{L,R},\bullet_{c},\circ_{s+1}\rangle\otimes|\alpha\rangle.$
@@
With $\alpha=-\lambda_{1}/(\sqrt{2}\omega)$. The braiding or exchange process between $\hat{\gamma}_{L}$ and $\hat{\gamma}_{R}$ MFs, as described by the operator $\hat{U}_{B}^{(L,R)}=e^{\frac{\pi}{4}\hat{\gamma}_{L}\hat{\gamma}_{R}}=\frac{1}{\sqrt{2}}\left(1+\hat{\gamma}_{L}\hat{\gamma}_{R}\right)$ (performing the exchange transformations $\hat{\gamma}_{L}\rightarrow\hat{\gamma}_{R}$ and $\hat{\gamma}_{R}\rightarrow-\hat{\gamma}_{L}$), when acting on the ground state produces the result $\hat{U}_{B}^{(L,R)}|{\rm GS}\rangle=e^{-i\frac{\pi}{4}}|{\rm GS}\rangle$. By using a local phase shifter acting just on the $j=s+1$ chain site [58, 59, 60, 46], a controlled complex coupling strength $\lambda_{2}e^{i\phi_{2}}$ should be possible enabling a cyclic evolution of the cavity coherent state $|\alpha\rangle$ and consequently to access to a specific Berry phase of the whole matter-photon state (dynamic phases which depend on the system’s specific time evolution may be experimentally canceled out [61, 62]).
@@
Now, we focus on recovering the same $e^{-i\frac{\pi}{4}}$ global phase factor on the GS by performing an adiabatic cyclic evolution of the coherent cavity field (Berry phase). The GS adiabatic evolution is fully contained in the two-dimensional sub-space spanned by the states $\{|{\rm GS}_{+}\rangle,|{\rm GS}_{-}\rangle\}$, as previously assured (for details see B). We perform a cyclic evolution of the GS by fixing a $\lambda_{2}>0$ value while varying the phase $\phi_{2}=\phi$ of the coupling strength, $0\leq\phi\leq 2\pi$. Thus, for the GS we have $|{\rm GS}(\phi)\rangle=C_{+}(\phi)|{\rm GS}_{+}(\phi)\rangle+C_{-}(\phi)|{\rm GS}_{-}\rangle$. Notice that the state $|{\rm GS}_{-}\rangle$ does not depend on the phase $\phi$. We proceed with an analytical calculation of the Berry phase using the sub-space-confined evolution of the GS. By direct calculation, we obtain:
$\varphi_{B}=\int_{0}^{2\pi}\frac{d\phi}{2\omega}\delta E_{\phi}\left[1-\frac{\delta E_{\phi}}{\sqrt{4\Delta^{2}|\langle\alpha_{-}|\alpha_{+}(\phi)\rangle|^{2}-\delta E^{2}_{\phi}}}\right].$ (13)
@@
The parity $P_{L,R}=i\langle\hat{\gamma}_{L}\hat{\gamma}_{R}\rangle$, between $\hat{\gamma}_{L}$ and $\hat{\gamma}_{R}$ MFs, is plotted in Fig. 4(b) as a function of $\lambda_{2}e^{i\phi_{2}}$, for $0\leq\lambda_{2}\leq\lambda_{1}$ and $0\leq\phi_{2}\leq 2\pi$. $P_{L,R}\rightarrow-1$ for $\lambda_{2}\approx 0$ in agreement with the GS given by Eq. (12). The solid red line shows the corresponding results for varying real $\lambda_{2}$ values from the positive to the negative region. At the same time, the closed curve depicts the $P_{L,R}$ evolution just at the $\lambda_{2}$ value producing the correct Berry phase of $\pi/4$, emulating the perfect braiding/exchange between $\hat{\gamma}_{L}$ and $\hat{\gamma}_{R}$ MFs. The $\pi/4$ phase emerges from the interplay of geometrical and topological features under these conditions, with its resilience tied to the robustness of the optical Berry phase readout.
Previous studies employing gate-controlled junctions between topological chains or nanowires have predicted braiding in two-dimensional scenarios, underscoring the need for more accurate modeling of strictly one-dimensional systems. We emphasize that in the present work we address this specific issue in a simpler 1D scenario, providing predictions based on the best available tools to model anyon statistics (fusion rules) and braiding in 1D KC-cavity systems.
5 Majorana-Schrödinger cat states